# Metric Pairs and Tuples in Theory and Applications

### v1

@@ Abstract.
${}^{\text{B}}$ M. Che has been funded in part by the Austrian Science Fund (FWF) [Grant DOI: 10.55776/STA32], and in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2047/1 – 390685813
${}^{\text{C}}$ M. Cuerno has been supported by the research grant PID2021-124195NB-C32 from the Ministerio de Economía y Competitividad de España (MINECO)
@@ 1. Introduction
The Gromov–Hausdorff distance and convergence have become fundamental tools for quantitatively comparing and studying metric spaces. Gromov introduced this concept in several papers [gromov_1, gromov_3, gromov_2] as a generalization of the classical Hausdorff distance, which compares closed subsets of the same metric space.
@@
Beyond its significant theoretical contributions in geometric group theory and differential geometry—such as Gromov’s compactness theorem [gromov_1, gromov_3, gromov_2]—the Gromov–Hausdorff distance has found numerous applications. In particular, Topological Data Analysis (TDA) uses it as a foundational tool in proving key stability theorems for objects like persistence diagrams (see [chazal2009gromov]). Moreover, it continues to serve as a reference for recent developments in stability and computational results, as seen in [adams_optimal, adams2024gromovhausdorffdistancesquotientmetric, adams2025hausdorffvsgromovhausdorffdistances, bubenik2018interleavinggromovhausdorffdistance, torrascasas2024stability].
@@
In [che_diagrams], the second named author and collaborators introduced a new notion of Gromov–Hausdorff convergence for metric pairs. Specifically, this Gromov–Hausdorff convergence for metric pairs extends the classical pointed Gromov–Hausdorff convergence (see, for example, [buragoburagoivanov]). Subsequently, the first and second named authors established the foundations of this theory for metric pairs and tuples, proving results such as embedding, completeness, and compactness theorems [aagmc]. Additionally, they presented a relative version of Fukaya’s theorem on quotient spaces under Gromov–Hausdorff equivariant convergence, as well as a version of the Grove–Petersen–Wu finiteness theorem for stratified spaces.
@@
Recent references to the formalism of metric pairs can be found in [alattar2024], where the author defined an equivariant version of the convergence of metric pairs, and in [sakovich-sormani2025], where the authors defined multiple notions of convergence of space-times, one of them referred to as the Future-Developed Intrinsic Distance, where the structure of metric pair arises when one considers space-times endowed with the so-called null distance and the distinguished subsets are Cauchy hypersurfaces.
@@
In this paper, we continue building the theory behind the Gromov–Hausdorff convergence and distance for metric pairs and tuples as well as presenting some applications that broaden the scope of this tool. First, after showing an equivalent formulation for the Gromov–Hausdorff distance for metric pairs in the sense of correspondences, we begin proving that this distance is geodesic and separable, as well as a generalization of the Arzelà–Ascoli theorem for metric pairs. The following theorems present some specific notation that will be introduced in sections 2 and 3.
@@
As an application of this framework, in section 4 we prove a version of the main theorem in [cassorla], stating the approximability of metric pairs by surfaces.
@@
Finally, to highlight the potential of the the Gromov–Hausdorff distances for metric pairs and tuples, in section 5 we show how these distances arise naturally in three applied settings: spaces of hypernetworks and graphs (see [needham_original, needham2024]); spaces of simplicial complexes (see [nnadi]); and persistence matching diagrams (see [torras_rocio_2, torras_rocio_1, torrascasas2024stability]).
@@ Acknowledgements.
The authors would like to thank Professor Christina Sormani, for fruitful conversations, and colleagues and friends for their invaluable comments at an early stage of this manuscript.
@@ 2.1. Gromov-Hausdorff distance
We present the basic definitions of the Hausdorff and Gromov–Hausdorff distance. For details and proofs of the following, we refer the reader to [buragoburagoivanov].
@@
Moreover, the hyperspace $H(X)$ of $X$, i.e., the space of closed subsets of a metric space $X$, endowed with the Hausdorff distance is a metric space.
@@ 2.2. Gromov-Hausdorff distance for metric pairs and tuples
The following notions are taken from [aagmc, che_diagrams].
@@
In the following definition, and in the sequel, we say that a metric pair $(X,A)$ is compact whenever the space $X$ itself is compact.
@@
where the infimum is taken over all admissible metrics $\delta$ in $X\sqcup Y$. We denote by $(\operatorname{GH}_{k},d_{\operatorname{GH}})$ the metric space of compact metric $k$-tuples (see theorem 2.13 of [aagmc]).
@@ 2.3. Equivalent formulation
[…]
@@ 3.1. Geodesicity
Lemma 3.2 (Extension of lemma 2.1 in [chowmemo]).
@@
$\left|\operatorname{dis}(R)-\operatorname{dis}(S)\right|$ $=\left|\frac{1}{2}\left(\sup_{(x,y),(x^{\prime},y^{\prime})\in R}\left|d_{X}(x,x^{\prime})-d_{Y}(y,y^{\prime})\right|\right.\right.$
@@
Let us take a sequence of non-negative numbers $\{\varepsilon_{n}\}_{n\in\mathbb{N}}$ which converges to 00. Now, for each $n$, let $A_{n}$ and $B_{n}$ be $\varepsilon_{n}/2$-nets for $A$ and $B$, respectively. We extend them to obtain $X_{n}$ and $Y_{n}$, $\varepsilon/2$-nets for $X$ and $Y$. By the definition of $\varepsilon$-net, we have that $d_{\mathrm{H}}^{\delta}((X_{n},A_{n}),(X,A))\leq\varepsilon$ and $d_{\mathrm{H}}^{\delta}((Y_{n},B_{n}),(Y,B))\leq\varepsilon$ for every admissible metric $\delta$. It implies that $(X_{n},A_{n})\to(X,A)$ and $(Y_{n},B_{n})\to(Y,B)$.
@@
Finally we have to prove that $R$ is actually a pair correspondence. For every $n\in\mathbb{N}$, using [chowmemo, Lemma 2.1] we get
@@
Using [chowmemo, Lemma 1.3], it is sufficient to prove
@@
[…]
@@ 3.2. Separability
$A^{\prime}=\{a^{\prime}\in X^{\prime}:(a,a^{\prime})\in R,\ \text{ for some }\ a\in A\}.$
@@ 3.3. Arzelà–Ascoli theorem for $\operatorname{GH}_{1}$
In this section we prove a version of the classical Arzelà–Ascoli theorem for metric pairs, generalizing the one presented in [herron2016] for maps between pointed metric spaces.
@@
$\operatorname{Lip}(f)=\inf\left\{K\in\mathbb{R}\cup\{\infty\}\left|\frac{d_{Y}(f(x_{1}),f(x_{2}))}{d_{X}(x_{1},x_{2})}\leq K\text{ for every }x_{1}\neq x_{2}\right.\right\},$
@@
By [aagmc, Theorem 1.1] we know that there are metric pairs $(\mathcal{X},\mathcal{A})$, $(\mathcal{Y},\mathcal{B})$ and isometric embeddings $\varphi_{i}\colon(X_{i},A_{i})\to(\mathcal{X},\mathcal{A})$ and $\psi_{i}\colon(Y_{i},B_{i})\to(\mathcal{Y},\mathcal{B})$ such that
@@
The rest of the proof follows along the same lines as that of [herron2016, Proposition 5.1]. The only detail to take into account is that the compactness of $A_{i}$ and the properness of $X_{i}$ for $i\in\mathbb{N}\cup\{\infty\}$ imply that the neighborhoods $\overline{B}_{R}(A_{i})$ ar compact for any $R>0$.
Finally, we obtain the last assertions from the proof of [herron2016, Proposition 5.1]. Namely, if $f_{i}$ (or their restrictions to $A_{i}$) are $L$-Lipschitz, let $u,v\in X_{\infty}$ (respectively, $u,v\in A_{\infty}$) and consider sequences $\{u_{i}\}_{i\in\mathbb{N}}$ and $\{v_{i}\}_{i\in\mathbb{N}}$ such that $u_{i},v_{i}\in X_{i}$ (or $u_{i},v_{i}\in A_{i}$, respectively), and $\varphi_{i}(u_{i})\to u$, $\psi_{i}(v_{i})\to v$. By the triangle inequality,
@@ 4. Approximation by surfaces
In [cassorla], Cassorla proved that any compact length space can be approximated by Riemannian surfaces. In this section, we prove a generalization of that result but for metric pairs.
@@
Recall that $D\subset X$ is a $\nu$-dense set if for each $x\in X$ there exists $a(x)\in D$ so that $d_{X}(x,a(x))<\nu$. Furthermore, $D$ is a minimal $\nu$-dense set if $D$ is $\nu$-dense and if $a_{1},a_{2}\in D$, $a_{1}\neq a_{2}$, then $d_{X}(a_{1},a_{2})>\nu$. We define the 1-complex $K=K(X,A,n)$ with vertices labeled by $A_{n}$ and 1-simplexes connecting $x_{i}$ to $x_{j}$ of length $d_{X}(x_{i},x_{j})$ if and only if
@@
Lemma 4.1 (Lemma 2.3 of [cassorla]).
@@
Lemma 4.2 (Example 2, section 2 of [petersen]).
@@
Now, the second step is to construct a surface near to the simplicial pair $(L,K)$. Let $0<\varepsilon<10^{-(n+1)}$, choose a minimal $\varepsilon$-dense set $\{y_{k}\}_{k=1}^{v}$ in $K$ and we extend it such that $\{y_{k}\}_{k=1}^{v}\subset\{y_{k}\}_{k=v+1}^{l}$ is a minimal $\varepsilon$-dense set of $L$ and these sets do not contain any vertices. Since $L$ is 1-dimensional and using Menger’s Einbettungssatz (Embedding’s theorem)(cf. page 285 of [menger]), we embed smoothly in $B_{\tilde{m}}(\bar{0})\subset\mathbb{R}^{3}$ with $\tilde{m}=\min_{i\neq j}\{d(x_{i},x_{j})\}$. Then we extend the image of the simplexes such that the length in $X$ is preserved.
@@
Part A bounds the section of the curve in the tube using equation (4.8), the distance in the circle $S_{f}$ and the minimality of $\{y_{j}\}_{i=1}^{l}$. And part B is a bound for the section on the sphere obtained with equation (4.5). Finally, A$+$B is smaller than $3\varepsilon$ due to the definition of $\bar{t}$ and $\bar{r}$.
@@
[…]
@@ 5.1. Hypernetworks
Recently, Needham and collaborators delved into the theory of hypergraphs and hypernetworks [needham_original, needham2024]. A hypergraph is a pair $(X,Y)$ where $X$ represents the set of nodes and $Y$ is a collection of subsets of $X$, called hyperedges. Additionally, an hypernetwork is a triple $(X,Y,\omega)$ such that $X$ and $Y$ are sets and $\omega\colon X\times Y\to\mathbb{R}$ is an arbitrary function. A hypergraph is an example of a hypernetwork where $\omega$ is the incidence function.
We can see a metric pair $(X,A)$ as a hypernetwork $(X\times A,X\times A,\omega_{X})$ with
@@
In [needham2024], a Gromov–Hausdorff distance for networks and hypernetwork is defined. In the context of metric pairs seen as hypernetworks, taking the metric pairs $(X,A)$ and $(Y,B)$ seen as hypernetworks, and a hypernetwork correspondence $R$ between $X\times A$ and $Y\times B$ we get:
$\operatorname{dis}_{\operatorname{net}}(R)$ $=$ $\sup\limits_{\begin{subarray}{c}(x,a,y,b)\in R\\ (x^{\prime},a^{\prime},y^{\prime},b^{\prime})\in R\end{subarray}}\left|\omega_{X}((x,a),(x^{\prime},a^{\prime}))-\omega_{Y}((y,b),(y^{\prime},b^{\prime}))\right|$
$=$ $\sup\limits_{\begin{subarray}{c}(x,a,y,b)\in R\\ (x^{\prime},a^{\prime},y^{\prime},b^{\prime})\in R\end{subarray}}\left|\frac{d_{X}(x,x^{\prime})+d_{X}(a,a^{\prime})-d_{Y}(y,y^{\prime})-d_{Y}(b,b^{\prime})}{2}\right|$
$\leq$ $\frac{1}{2}\left(\sup\limits_{\begin{subarray}{c}(x,a,y,b)\in R\\ (x^{\prime},a^{\prime},y^{\prime},b^{\prime})\in R\end{subarray}}\left|d_{X}(x,x^{\prime})-d_{Y}(y,y^{\prime})\right|+\sup\limits_{\begin{subarray}{c}(x,a,y,b)\in R\\ (x^{\prime},a^{\prime},y^{\prime},b^{\prime})\in R\end{subarray}}\left|d_{X}(a,a^{\prime})-d_{Y}(b,b^{\prime})\right|\right)$
@@
This means that $d_{\operatorname{N}}(X\times A,Y\times B)\leq d_{\operatorname{GH}}((X,A),(Y,B))$.

### v2

@@ 1. Introduction
The Gromov–Hausdorff distance and convergence are fundamental concepts in metric geometry, used to quantitatively compare and study metric spaces. Indeed, the Gromov–Hausdorff distance was introduced as a way to measure how close two different compact metric spaces are. Introduced by Edwards [undefs] and rediscovered by Gromov in a series of papers [undefab, undefad] (see [undefac] for an English translation), this concept is a generalization of the classical Hausdorff distance, which compares closed subsets of the same metric space.
The Gromov–Hausdorff framework has led to numerous important theoretical contributions. One of them is the celebrated Gromov’s compactness theorem [undefab], which has motivated fruitful theories of metric measure spaces that naturally arise as limits of sequences of Riemannian manifolds with lower curvature bounds in the Gromov–Hausdorff topology, such as Alexandrov spaces [undefh], Ricci limits [undefn, undefo, undefp] and, more generally, (R)CD spaces [undefd, undefy, undefaj, undefas, undefat]. See [undefc, undefw, undefx, undefaf, undefao, undefaq] for a glimpse of these developments. The Gromov–Hausdorff framework has also been used to formulate the stability of other geometric conditions [undefj, undefv, undefai].
@@
Beyond its theoretical significance, the Gromov–Hausdorff distance has found numerous applications. For example, Topological Data Analysis (TDA) uses it as a foundational tool in proving stability theorems for objects like persistence diagrams (see [undefk]; see also [undefe] for a brief introduction to TDA). Moreover, it continues to play a key role in recent work on stability theory and computational methods; see [undef, undeff, undefau, undefaw].
@@
In [undefl, undefm], Che and collaborators introduced a notion of Gromov–Hausdorff convergence for metric pairs, which extends the classical pointed Gromov–Hausdorff one (see, for example, [undefg, undefag]). Subsequently, Ahumada Gómez and Che established the foundations of this theory for metric pairs and tuples, proving results such as embedding, completeness, and compactness theorems [undefa]. Additionally, they presented a relative version of Fukaya’s theorem on quotient spaces under Gromov–Hausdorff equivariant convergence [undefv], as well as a version of the Grove–Petersen–Wu finiteness theorem [undefae] for stratified spaces.
@@
There are also recent references to the formalism of metric pairs. Alattar defined an equivariant version of the convergence of metric pairs in [undefb]. Additionally, Sakovich and Sormani introduced multiple notions of convergence of space-times, one of them referred to as the Future-Developed Intrinsic Distance, where the structure of metric pair arises when one considers space-times endowed with the so-called null distance and the distinguished subsets are Cauchy hypersurfaces [undefar].
There are other applied contexts where having a notion of distance between metric pairs and tuples seems to be useful. For instance, in spaces of hypernetworks and graphs, as defined in [undefr, undefam]. Also in spaces of simplicial complexes, as studied in [undefan]. Finally, within the theoretical development of TDA, notions analogous to the Gromov–Hausdorff distance considered in this work have been introduced for spaces of persistence diagrams [undefz, undefaa, undefau].
@@
In this paper, we continue building the theory of Gromov–Hausdorff convergence and distance for metric pairs and tuples as well as presenting some applications that broaden the scope of this tool. First, after showing an equivalent formulation for the Gromov–Hausdorff distance for metric pairs in the sense of correspondences, we begin proving that this distance is geodesic and separable, as well as a generalization of the Arzelà–Ascoli theorem for metric pairs. Together, these results provide a collection of key structural properties, demonstrating that the space of metric pairs has a well-behaved metric geometry suitable for approximation, convergence, and stability analyses. The following theorems present some specific notation that will be introduced in sections 2 and 3.
@@
As an application of this framework, in section 4 we prove a version of the main theorem in [undefi], stating the approximability of metric pairs by surfaces.
@@
Finally, to highlight the potential of the the Gromov–Hausdorff distances for metric pairs and tuples, in section 5 we compare the metrics between metric tuples with those that have been defined for spaces of hypernetworks and graphs [undefr, undefam]; spaces of simplicial complexes [undefan]; and spaces of persistence matching diagrams [undefz, undefaa, undefau].
@@ Acknowledgements.
The authors would like to thank Professor Christina Sormani, for fruitful conversations, and colleagues and friends for their invaluable comments at an early stage of this manuscript. A. Ahumada Gómez and M. Cuerno have been supported by the research grant PID2024-158664NB-C2 from Miciu/AEI/10.13039/501100011033/FEDER/UE. M. Cuerno has also been supported by the research grant PID2021-124195NB-C32 from the Ministerio de Economía y Competitividad de España (MINECO) and by the project “Charting political ideological landscapes in Europe: Fault lines and opportunities (POL-AXES)” - Programa Primas y Problemas de la Fundación BBVA 2023. M. Che was funded in part by the Austrian Science Fund (FWF) [Grant DOI: 10.55776/STA32], and in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2047/1 – 390685813. Part of this research was carried out at the Hausdorff Institute of Mathematics in Bonn, during the trimester program “Metric Analysis”. M. Che wishes to express his gratitude to this institution and the organizers of the program for the stimulating atmosphere and the excellent working conditions.
For open access purposes, the authors have applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission.
@@ 2.1. Gromov-Hausdorff distance
We present the basic definitions of the Hausdorff and Gromov–Hausdorff distance. For details and proofs of the following, we refer the reader to [undefg].
@@
Moreover, the hyperspace $H(X)$ of $X$, i.e., the space of closed subsets of a metric space $X$, endowed with the Hausdorff distance is a metric space whenever $X$ is compact.
@@ 2.2. Gromov-Hausdorff distance for metric pairs and tuples
The following notions are taken from [undefa, undefm].
@@
In the following definition, and in the sequel, we say that a metric pair $(X,A)$ is compact whenever the space $X$ itself is compact, which of course implies that $A$ is compact as well.
@@
where the infimum is taken over all admissible metrics $\delta$ in $X\sqcup Y$.
We denote by $(\operatorname{GH}_{k},d_{\operatorname{GH}})$ the metric space of (isometry classes of) compact metric $k$-tuples (see theorem 2.13 of [undefa]).
@@ 2.3. Equivalent formulation
Inequality ($\ast$) is satisfied because of the triangle inequality and ($\ast\ast$) is satisfied because $2r=\operatorname{dis}(R)$ and
@@ 3.1. Geodesicity
Lemma 3.2 (Extension of Lemma 2.1 in [undefq]).
@@
[…]
@@
Let us take a sequence of non-negative numbers $\{\varepsilon_{n}\}_{n\in\mathbb{N}}$ which converges to 0. Now, for each $n$, let $A_{n}$ and $B_{n}$ be $\varepsilon_{n}/2$-nets for $A$ and $B$, respectively. We extend them to obtain $X_{n}$ and $Y_{n}$, $\varepsilon/2$-nets for $X$ and $Y$. By the definition of $\varepsilon$-net, we have that $d_{\mathrm{H}}^{\delta}((X_{n},A_{n}),(X,A))\leq\varepsilon$ and $d_{\mathrm{H}}^{\delta}((Y_{n},B_{n}),(Y,B))\leq\varepsilon$ for every admissible metric $\delta$. It implies that $(X_{n},A_{n})\to(X,A)$ and $(Y_{n},B_{n})\to(Y,B)$.
@@
Finally we have to prove that $R$ is actually a pair correspondence. For every $n\in\mathbb{N}$, using [undefq, Lemma 2.1] we get
@@
Using [undefq, Lemma 1.3], it is sufficient to prove
@@
$=\left|t-s\right|\,\operatorname{dis}(R)=\left|t-s\right|\,2\,d_{\operatorname{GH}}((X,A),(Y,B)).$
@@ 3.2. Separability
[…]
@@ 3.3. Arzelà–Ascoli theorem for $\operatorname{GH}_{1}$
In this section we prove a version of the classical Arzelà–Ascoli theorem for metric pairs, generalizing the one presented in [undefag] for maps between pointed metric spaces.
@@
[…]
@@
By [undefa, Theorem 1.1] we know that there are metric pairs $(\mathcal{X},\mathcal{A})$, $(\mathcal{Y},\mathcal{B})$ and isometric embeddings $\varphi_{i}\colon(X_{i},A_{i})\to(\mathcal{X},\mathcal{A})$ and $\psi_{i}\colon(Y_{i},B_{i})\to(\mathcal{Y},\mathcal{B})$ such that
@@
The rest of the proof follows along the same lines as that of [undefag, Proposition 5.1]. The only detail to take into account is that the compactness of $A_{i}$ and the properness of $X_{i}$ for $i\in\mathbb{N}\cup\{\infty\}$ imply that the neighborhoods $\overline{B}_{R}(A_{i})$ ar compact for any $R>0$.
Finally, we obtain the last assertions from the proof of [undefag, Proposition 5.1]. Namely, if $f_{i}$ (or their restrictions to $A_{i}$) are $L$-Lipschitz, let $u,v\in X_{\infty}$ (respectively, $u,v\in A_{\infty}$) and consider sequences $\{u_{i}\}_{i\in\mathbb{N}}$ and $\{v_{i}\}_{i\in\mathbb{N}}$ such that $u_{i},v_{i}\in X_{i}$ (or $u_{i},v_{i}\in A_{i}$, respectively), and $\varphi_{i}(u_{i})\to u$, $\psi_{i}(v_{i})\to v$. By the triangle inequality,
@@ 4. Approximation by surfaces
In [undefi], Cassorla proved that any compact length space can be approximated by Riemannian surfaces. In this section, we prove a generalization of that result but for metric pairs.
@@
Recall that $D\subset X$ is a $\nu$-dense set if for each $x\in X$ there exists $a(x)\in D$ so that $d_{X}(x,a(x))<\nu$. The existence of a minimal $\nu$-dense set, for any $\nu>0$ follows from the compactness of $X$ (see, for example, [undefg, Exercise 1.6.4]).
Furthermore, $D$ is a minimal $\nu$-dense set if $D$ is $\nu$-dense and if $a_{1},a_{2}\in D$, $a_{1}\neq a_{2}$, then $d_{X}(a_{1},a_{2})>\nu$. We define the 1-complex $K=K(X,A,n)$ with vertices labeled by $A_{n}$ and 1-simplexes connecting $x_{i}$ to $x_{j}$ of length $d_{X}(x_{i},x_{j})$ if and only if
@@
Lemma 4.1 (Lemma 2.3 of [undefi]).
@@
Lemma 4.2 (Example 2, section 2 of [undefap]).
@@
Now, the second step is to construct a surface near to the simplicial pair $(L,K)$. Let $0<\varepsilon<10^{-(n+1)}$, choose a minimal $\varepsilon$-dense set $\{y_{k}\}_{k=1}^{v}$ in $K$ and we extend it such that $\{y_{k}\}_{k=1}^{v}\subset\{y_{k}\}_{k=v+1}^{l}$ is a minimal $\varepsilon$-dense set of $L$ and these sets do not contain any vertices. Since $L$ is 1-dimensional and using Menger’s Einbettungssatz (Embedding’s theorem)(cf. page 285 of [undefak]), we embed smoothly in $B_{\tilde{m}}(\bar{0})\subset\mathbb{R}^{3}$ with $\tilde{m}=\min_{i\neq j}\{d(x_{i},x_{j})\}$. Then we extend the image of the simplexes such that the length in $X$ is preserved.
@@
Part A bounds the section of the curve in the tube using equation (4.8), the distance in the circle $S_{f}$ and the minimality of $\{y_{j}\}_{i=1}^{l}$. And part B is a bound for the section on the sphere obtained with equation (4.5). Finally, A++B is smaller than $3\varepsilon$ due to the definition of $\bar{t}$ and $\bar{r}$.
@@
$\delta<2\varepsilon\text{ and }\{z_{i}\}_{i=1}^{l}\text{ is a $12\varepsilon$-dense set.}$
@@ 5.1. Hypernetworks
Recently, Needham and collaborators delved into the theory of hypergraphs and hypernetworks [undefr, undefam]. A network is a triple $(X,X,\omega)$, where $X$ is a set and $\omega\colon X\times X\to\mathbb{R}$ is an arbitrary function. A hypergraph is a pair $(X,Y)$ where $X$ represents the set of nodes and $Y$ is a collection of subsets of $X$, called hyperedges. More generally, an hypernetwork is a triple $(X,Y,\omega)$ such that $X$ and $Y$ are sets and $\omega\colon X\times Y\to\mathbb{R}$ is an arbitrary function. In particular, hypernetworks generalize networks, which arise as the special case $Y=X$ (see Example 2.8 of [undefam]). A hypergraph is an example of a hypernetwork where $\omega$ is the incidence function.
We can see a metric pair $(X,A)$ as a network $(X\times A,X\times A,\omega_{X})$ with
@@
In [undefam], a Gromov–Hausdorff distance for networks and hypernetwork is defined. In the context of metric pairs seen as networks, taking the metric pairs $(X,A)$ and $(Y,B)$ seen as networks, and a correspondence $R$ between $X\times A$ and $Y\times B$ we get:
$\operatorname{dis}_{\operatorname{net}}(R)$ $:=$ $\sup\limits_{\begin{subarray}{c}((x,a),(y,b)),\\ ((x^{\prime},a^{\prime}),(y^{\prime},b^{\prime}))\in R\end{subarray}}\left|\omega_{X}((x,a),(x^{\prime},a^{\prime}))-\omega_{Y}((y,b),(y^{\prime},b^{\prime}))\right|$
$=$ $\sup\limits_{\begin{subarray}{c}((x,a),(y,b)),\\ ((x^{\prime},a^{\prime}),(y^{\prime},b^{\prime}))\in R\end{subarray}}\left|\frac{d_{X}(x,x^{\prime})+d_{X}(a,a^{\prime})-d_{Y}(y,y^{\prime})-d_{Y}(b,b^{\prime})}{2}\right|$
$\leq$ $\frac{1}{2}\left(\sup\limits_{\begin{subarray}{c}((x,a),(y,b)),\\ ((x^{\prime},a^{\prime}),(y^{\prime},b^{\prime}))\in R\end{subarray}}\left|d_{X}(x,x^{\prime})-d_{Y}(y,y^{\prime})\right|+\sup\limits_{\begin{subarray}{c}((x,a),(y,b)),\\ ((x^{\prime},a^{\prime}),(y^{\prime},b^{\prime}))\in R\end{subarray}}\left|d_{X}(a,a^{\prime})-d_{Y}(b,b^{\prime})\right|\right)$
@@
This means that
$d_{\operatorname{net}}(X\times A,Y\times B):=\frac{1}{2}\inf\operatorname{dis}_{\operatorname{net}}(R)\leq d_{\operatorname{GH}}((X,A),(Y,B)),$
where the infimum runs over all correspondences $R$ between $X\times A$ and $Y\times B$.